OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 6) and (1, 3)?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ 6}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

just plug in your slope and the coordinates as given

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

but when I do that i get y= -3x+6 which what i should get after doing all that but the options for an answer is a. (y-3)= -3(x-1) b. (y+3) = 3(x+1) c. (y+3) = -3(x+1) d. (y-3) = 3(x+6)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@robtobey